Write the system of linear equations in trian-
gular form
x + y + z = 7
3x + 4y + z = 13
2x − 2y + 9z = 7
What is triangular form and how would I solve this problem using it?

Answer by stanbon(75887)

x + y + z = 7
3x + 4y + z = 13
2x − 2y + 9z = 7
----------------------
Subtract 3 times the 1st line from the 2nd line
Subtract 2 times the 1st line from the 3rd line
---
x + y + z = 7
0 + y - 2z = -8
0 -4y + 7z = -7
----------------------
Add 4 times the 2nd line to the 3rd to get:
x + y + z = 7
0 + y - 2z = -8
0 + 0 - z = -39
----
That is the triangle form (see the triangle of zeros in
the left bottom corner.
----
Note that you can solve for "z" in the 3rd line z = 39
--
Then you can substiute 39 for "z" in the 2nd line and solve for y = 70.
---
Then you can substitute for z and for y in the 1st line and solve for x=-102.
======================
Cheers,
Stan H.